My answer does not work-I am not sure what I did wrong!
The length of a rectangle is 3 ft more than twice its width. The area of the rectangle is 35ft^2. Find the dimensions.
Thanks for your help!

Answer by mananth(16946)

width = w
length = 2w+3
Area = L*W
w(2w+3)=35
2w^2+3w=35
2w^2+3w-35=0
2w^2+10w-7w-35=0
2w(w+5)-7(w+5)=0
(w+5)(2w-7)=0
w=-5 OR w=7/2
ignore negative value
width = 3.5
length =>2w+3 => 2*3.5 +3 = 10